Solve
(3x+2)/4=(9)/(4)-(3-2x)/(6)
Please I need help with this. Thank you!

Answer by stanbon(75887)

Solve
(3x+2)/4=(9)/(4)-(3-2x)/(6)
-----
Multiply thru by 12 to get:
3(3x+2) = 3*9 - 2(3-2x)
-------
9x + 6 = 27 - 6 + 4x
5x = 27-12
5x = 15
x = 3
============
Cheers,
Stan H.
============